OpenStudy (anonymous):

Find y as a function of x if y(4)−10y′′′+25y′′=0, y(0)=15, y′(0)=1, y′′(0)=25, y′′′(0)=0.

OpenStudy (anonymous):

the y(4) means 4th derivative btw

zepdrix (zepdrix):

4th degree linear? So umm, the characteristic equation will be given by:\[\Large r^4-10r^3+25r^2\quad=\quad 0\]Factoring:\[\Large r^2\left(r^2-10r+25\right)\quad=\quad0\]Which further factors,\[\Large r^2(r-5)^2\quad=\quad 0\]

zepdrix (zepdrix):

Hmmm so it looks like we have 2 sets of repeated roots. \(\Large r=0, \qquad r=5\)

zepdrix (zepdrix):

For repeated roots of this form, we multiply by x each time the root repeats if I remember correctly. So our general solution would look something like:\[\Large y\quad=\quad c_1e^0+c_2xe^0+c_3e^5+c_4xe^5\]Which we could probably simplify a lil bit.\[\Large y\quad=\quad c_1+c_2x+(c_3+c_4x)e^5\]

zepdrix (zepdrix):

And then we've got a bunch of initial data to deal with? Oh boy :b

OpenStudy (unklerhaukus):

\[ y^{(4)}−10y'''+25y''=0,\quad y(0)=15,\quad y'(0)=1,\quad y''(0)=25,\quad y'''(0)=0\] \[ \mathcal L\Big\{y^{(4)}−10y'''+25y''\Big\}=\mathcal L\Big\{0\Big\}\]

OpenStudy (unklerhaukus):

\[\mathcal L\Big\{y^{(4)}\Big\}−10\mathcal L\Big\{y'''\Big\}+25\mathcal L\Big\{y''\Big\}=\mathcal L\Big\{0\Big\}\] \[\Big(s^4Y-s^3y(0)-s^2y'(0)-sy''(0)-y'''(0)\Big)\\−10\Big(s^3Y-s^2y(0)-sy'(0)-y''(0)\Big)\\+25\Big(s^2Y-sy(0)-y'(0)\Big)=0\]